# A Simple and Efficient Joint Measurement Strategy for Estimating Fermionic Observables and Hamiltonians

### v2

@@ IV.1 Proof for Joint Measurement strategy
Below we provide proof of Proposition 1:
@@ IV.2 Implementation of measurement strategies
Figure 4: Two exemplary realizations of the Jordan-Wigner encoding in systems of qubits that exhibit different spatial locality: a) 1D array and b) $L/2\times(L+1)$ rectangular qubit array (for $L=4$). The individual qubits are indicated by consecutive pairs within the blue rounded rectangles labelled by $Q_{1},Q_{2},\ldots$, each enclosing two Majorana mode operators indicated by black dots. Importantly, locality in Majorana modes (e.g. coupling $\mathrm{exp}(\mathrm{i}\varphi\gamma_{i,j})$ for neighboring Majorana modes $i,j$) does not in general correspond to locality with respect to the structure of the qubit lattice.
@@
It is more challenging to implement the unitaries $U_{\mathrm{pair}}$ and $U_{\mathrm{resh}}^{(r)}$. Note that, as explained in Sec. II.3, they are purely vertical (see (12)). That is, the permutations $\pi$ and $\sigma^{(r)}$ that correspond to $U_{\mathrm{pair}}$ and $U^{(r)}_{\mathrm{resh}}$ respectively, correspond to orthogonal transformations $R_{\pi/\sigma^{(r)}}=\oplus_{\beta=1}^{L/2}\tau^{\beta}$, where $\tau^{\beta}$ only permutes elements in the $\beta$–th column of qubits in the 2D layout under JW encoding. Importantly, in [52] (see also [55]), it was shown that such a class of FLO gates can be implemented in depth $\mathcal{O}(L)=\mathcal{O}(N^{1/2})$ and using $\mathcal{O}(L^{3})=\mathcal{O}(N^{3/2})$ two-qubit gates, by utilizing one extra auxiliary qubit per row. In Appendix G we provide the proof that these scalings hold for general FLO transformations, extending the treatment of number preserving FLO transformations in [52]). Thus, $U_{\mathrm{pair}}$ and $U^{(r)}_{\mathrm{resh}}$ can also be implemented with the claimed gate depth and count.
@@ Acknowledgements
We would like to thank Ingo Roth, Martin Kliesch and Alexander Gresch for helpful discussions, as well as Katarzyna Kowalczyk-Murynka for assistance with numerical computations. The authors acknowledge financial support from the TEAM-NET project co-financed by the EU within the Smart Growth Operational Program (contract no. POIR.04.04.00-00-17C1/18-00). D.M. acknowledges support from PNRR MUR Project No. PE0000023-NQSTI. The authors declare that they have no competing interests.
Appendix A Number of measurement settings ensuring all quadruples are measured in 2D setting
@@
Appendix B Four measurement rounds are sufficient to estimate Hamiltonians
@@
Appendix C Details of the strategy for physical fermionic Hamiltonians
@@
Appendix D Variance of the Hamiltonian estimator
@@
Appendix E Optimization of coefficients in estimator for physical fermionic Hamiltonians
@@
Table 2: Comparison of variances for estimating energies of standard benchmark molecules using the joint measurability strategy, with coefficients $\alpha^{(r)}_{A}$ in Eq. (14) either uniform (32) or optimized (28) for the ground state of the Hamiltonian. The variances are calculated for the ground states of five molecular Hamiltonians, with units Ha2 (see also Table 1 of the main text which summarizes only the optimized coefficients). Except for H2, in which both lower-flat [43] and almost Hadamard [56] matrices are used, we typically apply the lower-flat orthogonal matrices defined in Appendix B.1. of [36].
@@
Appendix F Details of the numerical benchmarks on molecular Hamiltonians
@@
Appendix G Shallow circuits for vertical FLO unitaries in 2D layouts under the JW encoding
@@
In this appendix we will show that a class of unitaries we call JW-vertical FLO transformations, which include $U_{\mathrm{pair}}$ and $U^{(r)}_{\mathrm{resh}}$, can be implemented in the horizontal JW encoding with the required gate depth and gate count. The basic idea, as demonstrated in the proof of Prop. 3, is to find a suitable map between the relevant circuit in the vertical JW encoding (for which the implementation is trivial) and its realization in the horizontal JW encoding.
G.1 Majorana mode labellings
Figure 5: Two different ways of labelling $2N=2l_{x}l_{y}$ Majorana modes in a $l_{x}\times l_{y}$ rectangular qubit lattice for $l_{x}=2,l_{y}=5$. Black dots represent Majorana modes while qubits are represented by blue rounded rectangles. Part a) is the standard labelling (left to right in each row) used througout the main text. Part b) is a snake-like labelling used in [52, 55] and in this appendix.
@@ G.1 Majorana mode labellings
G.2 Qubit labellings, encodings and transformations
Figure 6: Example of qubit labellings, encodings and transformations used in this appendix, for $2N=2l_{x}l_{y}$ Majorana modes encoded in $l_{x}\times l_{y}$ rectangular qubit lattice for $l_{x}=2,l_{y}=5$. Dots in the figures represent Majorana modes while qubits are represented by blue rounded rectangles. Parts a) and b) present horizontal JW and vertical JW qubit encodings, respectively. These encodings differ by the locality structure of the encoded Majorana modes, indicated by the green "snake". In the horizontal (vertical) encoding, neighboring modes are arranged along a "snake" which traverses subsequent qubit rows horizontally (columns vertically). Part c) represents a snake labelling of qubits ($Q_{1}-Q_{10}$), as well as two JW-vertical (see (33) for definition) transformations $V_{I},V_{II}$, i.e. transformations which only act on modes contained in one column of qubits.
@@ G.2 Qubit labellings, encodings and transformations
G.3 Compilation of vertical FLOs under horizontal JW encoding
@@
Figure 7: A graphical presentation of a circuit implementing a unitary $\Gamma$ for $l_{x}=3$, $l_{y}=3$. Black and blue dots represent the system and ancillary qubits respectively. The subsequent diagrams depict subsequent stages of the computation. A detailed explanation of different parts of the circuit is given in the last part of the proof of Proposition 3. Specific steps in the figure are referenced in text by their row (I-V) and column (A-D). The qubits $i-j$, whose parities are stored in a given ancillary qubit, are indicated by annotation P($i-j$) placed below/above the ancillary qubit.

### v3

@@ IV.1 Proof for joint measurement strategy
Below we provide the proof of Proposition 1:
@@ IV.2 Implementation of measurement strategies
Figure 4: Two exemplary realizations of the Jordan-Wigner encoding in systems of qubits that exhibit different spatial locality.
a) 1D array and b) $L/2\times(L+1)$ rectangular qubit array (for $L=4$). The individual qubits are indicated by consecutive pairs within the blue rounded rectangles labelled by $Q_{1},Q_{2},\ldots$, each enclosing two Majorana mode operators indicated by black dots. Importantly, locality in Majorana modes (e.g. coupling $\mathrm{exp}(\mathrm{i}\varphi\gamma_{i,j})$ for neighboring Majorana modes $i,j$) does not in general correspond to locality with respect to the structure of the qubit lattice.
@@
It is more challenging to implement the unitaries $U_{\mathrm{pair}}$ and $U_{\mathrm{resh}}^{(r)}$. Note that, as explained in Sec. II.3, they are purely vertical (see (12)). That is, the permutations $\pi$ and $\sigma^{(r)}$ that correspond to $U_{\mathrm{pair}}$ and $U^{(r)}_{\mathrm{resh}}$ respectively, correspond to orthogonal transformations $R_{\pi/\sigma^{(r)}}=\oplus_{\beta=1}^{L/2}\tau^{\beta}$, where $\tau^{\beta}$ only permutes elements in the $\beta$–th column of qubits in the 2D layout under JW encoding. Importantly, in [52] (see also [55]), it was shown that such a class of FLO gates can be implemented in depth $\mathcal{O}(L)=\mathcal{O}(N^{1/2})$ and using $\mathcal{O}(L^{3})=\mathcal{O}(N^{3/2})$ two-qubit gates, by utilizing one extra auxiliary qubit per row. In Supplementary Information H we provide the proof that these scalings hold for general FLO transformations, extending the treatment of number preserving FLO transformations in [52]). Thus, $U_{\mathrm{pair}}$ and $U^{(r)}_{\mathrm{resh}}$ can also be implemented with the claimed gate depth and count. Numerical experiments in [55], which consider the complexity of the method for passive FLO transformations, adapted from [52], suggest that the gate scaling advantage becomes apparent for systems larger than 18$\times$18 qubits due to the gate overhead incurred by the method of implementing the vertical transformations. We therefore expect, in practice, that our method exhibits advantageous gate scaling in the 2D setting for system sizes of the order of hundreds of qubits.
Data Availability
The atomic coordinates of the molecular Hamiltonians used for numerical benchmarks are available in the Supplementary Information F. The code utilized for the benchmarks is available upon request.
@@ Acknowledgements
We would like to thank Ingo Roth, Martin Kliesch and Alexander Gresch for helpful discussions, as well as Katarzyna Kowalczyk-Murynka for assistance with numerical computations. The authors acknowledge financial support from the TEAM-NET project co-financed by the EU within the Smart Growth Operational Program (contract no. POIR.04.04.00-00-17C1/18-00). D.M. acknowledges support from PNRR MUR Project No. PE0000023-NQSTI. M.O. acknowledges support from QuantERA Project TouQan financed by the National Science Center No. 2023/05/Y/ST2/00140 (ERA-NET Cofund QuantERA II).
Author contributions
M.O. developed the initial research idea and oversaw all aspects of the project. D.M. and M.O. were involved in the construction of the joint measurement and its application to estimating Majorana pairs, quadruples and Hamiltonians. J.M., D.M. and M.O. participated in developing concrete sample efficient schemes with favourable gate depths and counts. J.M. conducted all numerical calculations. All authors contributed to the discussions, analytical calculations, and writing of the manuscript.
Supplementary Information A Number of measurement settings ensuring all quadruples are measured in 2D setting
@@
Supplementary Information B Four measurement rounds are sufficient to estimate Hamiltonians
@@
Supplementary Information C Details of the strategy for physical fermionic Hamiltonians
@@
Supplementary Information D Variance of the Hamiltonian estimator
@@
Supplementary Information E Optimization of coefficients in estimator for physical fermionic Hamiltonians
@@
Table 2: Comparison of variances for estimating energies of standard benchmark molecules using the joint measurability strategy.
The coefficients $\alpha^{(r)}_{A}$ in Eq. (14) are either uniform (32) or optimized (28) for the ground state of the Hamiltonian. The variances are calculated for the ground states of five molecular Hamiltonians, with units Ha2 (see also Table 1 of the main text which summarizes only the optimized coefficients). Except for H2, in which both lower-flat [44] and almost Hadamard [56] matrices are used, we typically apply the lower-flat orthogonal matrices defined in Appendix B.1. of [37].
@@
Supplementary Information F Details of the numerical benchmarks on molecular Hamiltonians
@@
Supplementary Information G Details on the numerical simulation of the scheme for fermionic Gaussian states
Data: number of samples $s$;
covariance matrix $C_{\rho}$ of Gaussian state $\rho$;
orthogonal matrices $R_{U}^{(r)}$ ($r\in[4]$) defining FLO transformations $U^{(r)}$ of the rounds from Section II.4.
Result: $s\times 4$ samples $\{\bm{q}_{j}^{(r)}\}\,(j\in[s],r\in[4]$), $\bm{q}=(q_{1},q_{2},\ldots,q_{N})$, $q_{i}=\pm 1$ ;
  $\triangleright$ (step (iii) from Section II.2)
bit-strings $\{X^{(r)}_{j}\}\,(j\in[s],r\in[4])$ corresponding to random Majoranas drawn for each sample.
for $j\leftarrow 1$ to $s$ do
       for $r\leftarrow 1$ to $4$ do
             $C\leftarrow C_{\rho}$ ;
              $\triangleright$ initialize a covariance matrix of the input state
            draw random subset $X\subseteq 2N$ defining Majorana $\gamma_{X}$;
            $X^{(r)}_{j}\leftarrow X$;
            construct $R_{X}$ with matrix elements $R_{X,kl}=-1^{|X|+|X\cap\{k\}|}\delta_{kl}$, s.t. $\forall_{k\in[2N]}$ $\gamma_{X}\gamma_{k}\gamma_{X}^{\dagger}=\sum_{l}R_{X,kl}\gamma_{l}$;
            $C\leftarrow R_{U}^{(r)T}R_{X}^{T}CR_{X}R_{U}^{(r)}$ ;
              $\triangleright$ evolve the state with $\gamma_{X}^{\dagger}$ and $U^{\dagger}$
            $\bm{x}\leftarrow\mathrm{FLOYao.sample}(C)$ ($\bm{x}=(x_{1},x_{2},....,x_{N}),x\in\{0,1\}$) ;
              $\triangleright$ measure in computational basis
            $q_{i}\leftarrow 1-2x_{i}$, where $\bm{q}_{j}^{(r)}=(q_{1},q_{2},...,q_{N});$
       end for
end for
Algorithm 1 Classical simulation of the sampling procedure for estimating physical Hamiltonians (Section II.4)
Data: $s\times 4$ samples $\{\bm{q}_{j}^{(r)}\}\,(j\in[s],r\in[4]$), $\bm{q}=(q_{1},q_{2},\ldots,q_{N})$, $q_{i}=\pm 1$;
orthogonal matrices $R_{U}^{(r)}$ ($r\in[4]$) defining FLO transformations $U^{(r)}$ of the rounds from Section II.4;
bit-strings $\{X^{(r)}_{j}\}\,(j\in[s],r\in[4])$ corresponding to random Majoranas drawn for each sample;
Hamiltonian $H=\sum_{A\in\mathcal{X}_{2}\cup\mathcal{X}_{4}}h_{A}\gamma_{A}$;
coefficients $\bm{\alpha}^{(r)}=(\alpha^{(r)}_{A_{1}},\alpha^{(r)}_{A_{2}},...,\alpha^{(r)}_{A_{h}})$ ;
  $\triangleright$ uniform or optimized, see Supplementary Information E
Result: Empirical estimator of Hamiltonian $\hat{\bar{H}}$ and variance $\mathrm{Var}(\hat{\bar{H}})$.
for $j\leftarrow 1$ to $s$ do
       for $r\leftarrow 1$ to $4$ do
            for $A\in\mathcal{X}_{2}\cup\mathcal{X}_{4}$ do
                  if $A\in\mathcal{M}^{(r)}$ then
                         choose $B^{\prime}\subset[N],B\subset[2N]$ according to Supplementary Information C;
                        $\eta_{A}^{(r)}\leftarrow{\mathrm{det}(R_{U,AB}^{(r)})}$ ;
                          $\triangleright$ sharpness
                        $e_{A}\leftarrow\mathrm{sgn}(\mathrm{det}(R_{U,AB}))(-1)^{|A\cap X|}\prod_{k\in B^{\prime}}q^{(r)}_{j,k}$ ;
                          $\triangleright$ step (iv) from Section II.2
                        $\hat{\gamma}_{A}^{(r)}\leftarrow\frac{1}{\eta_{A}^{(r)}}e_{A}$;
                   end if
                  else
                        $\hat{\gamma}_{A}^{(r)}\leftarrow 0$ ;
                          $\triangleright$ $\gamma_{A}$ not estimated from this round
                   end if
             end for
            $H^{(r)}\leftarrow\sum_{A}\alpha_{A}^{(r)}h_{A}\hat{\gamma}_{A}^{(r)}$;
       end for
      $H_{j}\leftarrow\sum_{r=1}^{4}H^{(r)}$;
end for
$\hat{\bar{H}}\leftarrow\frac{1}{s}\sum_{j=1}^{s}H_{j}$;
$\mathrm{Var}(\hat{\bar{H}})\leftarrow\frac{1}{s-1}\sum_{j=1}^{s}(H_{j}-\hat{\bar{H}})^{2}$;
Algorithm 2 Postprocessing of samples from Algorithm 1 for estimation of Hamiltonian
We simulate the strategy for physical Hamiltonians from Section II.4 for random fermionic Gaussian states, which is classically efficient [39, 46], and we numerically validate the accuracy of the variance formula (25) for Hamiltonian estimation. The simulation yields energy estimates that agree with analytical values, up to statistical precision. For each molecular Hamiltonian we draw a different random Gaussian state and simulate our joint measurement strategy using $\mathcal{O}(10^{5})$ rounds. The energies of each state are of the order of 1, and the values from analytical calculations and numerical estimations are in agreement up to at least one decimal place. This is orders of magnitude smaller than the variances, which are of orders comparable to the ground states in Table 1 ($10-10^{3}$).
Above in Algorithm 1 we present the procedure of obtaining the samples for the measurement scheme from classical simulation and in Algorithm 2 the postprocessing of the samples for Hamiltonian estimation. For the classical simulation we use the package FLOYao.jl [48] (see documentation of the package and [47] for an accessible introduction to the topic).
Supplementary Information H Shallow circuits for vertical FLO unitaries in 2D layouts under the JW encoding
@@
In this section we will show that a class of unitaries we call JW-vertical FLO transformations, which include $U_{\mathrm{pair}}$ and $U^{(r)}_{\mathrm{resh}}$, can be implemented in the horizontal JW encoding with the required gate depth and gate count. The basic idea, as demonstrated in the proof of Prop. 3, is to find a suitable map between the relevant circuit in the vertical JW encoding (for which the implementation is trivial) and its realization in the horizontal JW encoding.
H.1 Majorana mode labellings
Figure 5: Two different ways of labelling $2N=2l_{x}l_{y}$ Majorana modes in a $l_{x}\times l_{y}$ rectangular qubit lattice for $l_{x}=2,l_{y}=5$.
Black dots represent Majorana modes while qubits are represented by blue rounded rectangles. Part a) is the standard labelling (left to right in each row) used througout the main text. Part b) is a snake-like labelling used in [52, 55] and in this section.
@@ H.1 Majorana mode labellings
H.2 Qubit labellings, encodings and transformations
Figure 6: Example of qubit labellings, encodings and transformations used in this section, for $2N=2l_{x}l_{y}$ Majorana modes encoded in $l_{x}\times l_{y}$ rectangular qubit lattice for $l_{x}=2,l_{y}=5$.
Dots in the figures represent Majorana modes while qubits are represented by blue rounded rectangles. Parts a) and b) present horizontal JW and vertical JW qubit encodings, respectively. These encodings differ by the locality structure of the encoded Majorana modes, indicated by the green "snake". In the horizontal (vertical) encoding, neighboring modes are arranged along a "snake" which traverses subsequent qubit rows horizontally (columns vertically). Part c) represents a snake labelling of qubits ($Q_{1}-Q_{10}$), as well as two JW-vertical (see (33) for definition) transformations $V_{I},V_{II}$, i.e. transformations which only act on modes contained in one column of qubits.
@@ H.2 Qubit labellings, encodings and transformations
H.3 Compilation of vertical FLOs under horizontal JW encoding
@@
Figure 7: A graphical presentation of a circuit implementing a unitary $\Gamma$ for $l_{x}=3$, $l_{y}=3$.
Black and blue dots represent the system and ancillary qubits respectively. The subsequent diagrams depict subsequent stages of the computation. A detailed explanation of different parts of the circuit is given in the last part of the proof of Proposition 3. Specific steps in the figure are referenced in text by their row (I-V) and column (A-D). The qubits $i-j$, whose parities are stored in a given ancillary qubit, are indicated by annotation P($i-j$) placed below/above the ancillary qubit.